A traffic control engineer reports that 75% of the vehicles passing through a checkpoint are from within the state. What is the probability that fewer than 2 of thenext 5 vehicles are from out of state?
Answer #1
Given X~Binomial(n=5, p=0.25)
P(X=x)=xC5*(0.25^x)*(0.75^(5-x))

So the probability is
P(X<2)=P(X=0)+P(X=1)
=0C5*(0.25^0)*(0.75^(5-0))+1C5*(0.25^1)*(0.75^(5-1))
=0.6328125
answered by: dennis